Hi all, can anyone show me how to solve this equation;
Solve for x: y = x^3 - 4
Any help would be great.
Thanks, -Nick.

Answer by stanbon(75887)

Solve for x: y = x^3 - 4
-------------
y+4 = x^3
x = (y+4)^(1/3)
======================
Cheers,
Stan H.